18th May 2009 APM Meeting tables 11:30-13:00 Present: MJI, STH, SC, RWA, NAW, MR, EGS, RGM, JRL(later) Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Recent and upcoming meetings 4. Data archives update - AAT, ING, WFCAM, UKIRT 5. Optical/NIR processing - HAWKI, INT WFC, MegaCam, Subaru 6. WFCAM update 7. VISTA update 8. AOB - rolling grant status etc.. Minutes ------- 1. Actions from the last meeting STH RWA still to check PLONE pages for poorly aligned text. We need to populate the RWA Vista pages (JRL). JRL JRL HAWK-I web page: stalled. JRL has been processing data for RGM - specifically narrow-band photometry of galaxy clusters. MJI Stage I software release - still ongoing. MJI suggested we put some of JRL's catalogue software on it and EGS would add his program for converting binary tables as well. MJI Technical page - ongoing. MJI Utility to convert binary tables still ongoing. EGS UVEX support - program written but not yet sent. NAW VIZIER - IPHAS. NAW said he thought this might better be done verbally, in any case it could be safely left until August. In the meantime it is ongoing optimally slowly. STH ASTEP ongoing. MR Fits header for Mar 12. No reply about this from Andy Adanmson. MJI says we make up the header and drop this action. MJI Vista processing done. MJI Still waiting for the VISTA I/O disks. To be done next week after the A/C is fixed. JRL ESO archive: They have no intention of using Blu-Ray with us. JRL Tile headers: I've sent a night's data to ES at WFAU for them to ingest and hopefully will get some feedback on what they feel is missing. Although there is no tiling in the pipeline yet, I don't think there are any additional header items that will be generated when there are. SC Wall mounted monitor on lecture theatre wall costs £1800 for 1920 x 1080 pixels. Better and (and cheaper) to go for a 24-inch computer terminal (£300) or 42-inch TV (£500)? NAW suggested taking a laptop to Curry's and plugging it in. 2. WFAU minutes DR6 is in preparation. The data to the end of 08B is being held up at IoA due to lack of necessary data required to fix the Petrosian problem. Steve Warren wants the jpegs(?) checked by June 21. Action: MJI to find out the release date of DR6. At the UKIRT board meeting MJI was shown a new server room and lots of shiny new kit. RGM pointed out that extinction was being applied in the wrong sense and was now fixed. He also added something about fixed list-driven photometry - didn't catch a lot of this but MJI is actioned to fix it. 3. Recent and upcoming meetings MJI gave a talk at the WFCAM board meeting on May 22 and pointed out the problems with lack of personnel and funding. Ths was noted by Simon Berry and `they' are fully aware of the situation. Not talked to ESO directly - its politically sensitive between ESO and STFC at the moment due to penalty clauses on VISTA. An accelerated plan for LAS was presented at the meeting. LAS was due to finish at the end of 09B but it won't due to bad weather. STH said that there were big overheads with LAS due to lack of guide stars. VISTA IOT telecon was rescheduled to the end of June and then cancelled. Minutes of the previous telecon are available on the web pages. NAW gave a talk about PLATO data flow including the integration of ground-based follow-up of science data. He is confident it will continue with the launch expected at the end of 2017. There is an galactic plane consortium for EGAPS in September and NAW, MJI, EGS and maybe JRL will go. UKIRT workshop at ROE from Sept 14-16 with the deadline for submissions in mid-July. STH has been invited to talk about WFCAM by Ian Robson. JRL has registered for ADASS. We should put together an abstract on VISTA, which is due July 15. (Action: JRL) NAW said there is a GAIA Great meeting in Nice from Nov 19-20 and another here on Jul 7-8. This will be a small WG meeting about wide-field spectroscopy. 4. Data archives AAT: Data now transferred daily. ING and WFCAM also coming in regularly. JRL is still investigating UKIRT tapes that won't read. AA has said that we could FTP any missing datasets from UKIRT, so we'll compile a list of missing datasets in the next week or so and submit them to him. Hopefully it won't entail too much work for JAC. MR is automating the WFCAM reduction - he has written CRON jobs which get everything ready by 0830. The data arrives within 24 hours of being taken and MJI and MR take turns to reduce it. 5. Optical/NIR processing : HAWKI, INT WFC, MegaCam, Subaru HAWK-I: JRL has reduced two more datasets for RGM - a PI program of narrow-band photometry of galaxy clusters. Re-photometry has been done by MJI who has received complaints for his pains because he has not applied corrections for which we do not have the data! STH reports no further progress with zoomifying. INT WFC data is being processed for Johan Knapen with more to do. MJI is reducing some data for Vasily Belokurov. MJI says he plans to do some nebulosity tests in the galactic plane for IPHAS. EGS has reduced a few service nights for Alberto Fernandez Soto - UVRHalpha imaging and spectroscopy of elliptical galaxies to identify satellites. EGS to be involved as a co-author. H alpha filters for VST have been received in Munich and will be tested by Robert Greimel. 6. WFCAM update. Jim says there have been no notable changes to the pipeline. MJI says processing of 09A data is on schedule. Two nights were lost when there was a problem with the top-end. but which has now been fixed. We hve not received any non-linear sequences at low light levels yet, there is a problem with the dome illumination. WFAU transfers are up to date. MJI in touch with WFAU about re-processed data. a) GPS nebulosity fields - a while back. b) UDS data requiring a lot of extras before it can be ingested at ROE. Needs a unique version number in the FITS header. c) Sky files have the same name so they need to be re-named and the FITS headers updated. Apply photometric calibration with detector update. MJI is waiting to hear that this is completed then he will overwrite (the data, headers?). Petrosian? MJI is waiting to re-process the Orion date - from prime focus for Casali and ? There has been some discussion about filter fluxes on WFCAM. The filters are used in the survets and this measn a significant overhead . Someonewho?) wiull look at this. 7. VISTA update RWA said data is now coming from Garching on blackberry-sized 500 GByte disks, but still in oversized boxes. He has contacted Garching to see if the volume (and hence transport cost) can be reduced. STH spent 3 weeks at VISTA contributing to the commissioning effort. Noteworthy measurements made include (but not limited to): + Linearity Sequences + Flatfields + Very deep J-band field (total of 16 hours) + A number of SIQ (system image quality) tests (although seeing during STH's run was never perfect) + Lots of touchstone fields + A touchstone field stepped across the array to test for an illumination correction + The VISTA throughput (which is extremely good), measured values listed here (using zero points derived by MJI from data taken late last year before the Z-filter was available): Y 0.48, J 0.55, H 0.71, K 0.67 + A few tiles in a variety of regions - this will enable us to test our tiling software. + Use of the VISTA DIMM and ESO DIMM archive to measure the site seeing but as these figures were often larger than the Vista value it led to problems when subtracting seeing. STH has persuaded JPE to do most of the reduction here. STH's plans are to focus initially on the SIQ across the entire array for data taken in the very best conditions, taking account of the site seeing. JRL sent the following report: I have continued to make bug fixes to the summit pipeline as I find them. The science pipeline is now almost complete and is being debugged. There are still some notable omissions, (no tiling yet, WCS isn't yet back propagated to lower level data products) but I feel I can get this sorted by the end of June. There are still tickets in the ESO system regarding the summit pipeline that I have not yet addressed. These are mostly comments and don't require any update to the software. Two data discs with approximately a fortnight's data have arrived. The second only arrived yesterday so I haven't had a chance to look at it yet. Most of the data for the first disc concerns verification tests being done at the telescope. I did an initial reduction of the first weeks data as though they were science nights. MJI has discovered a few major problems which have now been rectified and I will re-reduce those nights this weekend. Once that is done MJI can have another look. We have done a great deal of work to help with the verification tests for VISTA: Linearity: Linearity sequences were taken in all the filters. These were analysed to create linearity curves for each detector for all filters. This turned up a serious problem with a lack of stability in the dome lights. One of the linearity curves worked out reasonably well and will be used for the reduction of the current data. Read noise and Gain: Data from the linearity sequences were reused to work out estimates of gain and readout noise for each of the detectors. The lack of stability of the dome light source does make these estimates a little uncertain. In order to test for excess read noise induced by the telescope, dome tracking, autoguider etc, dark observations were done of varying exposure times during the day when most of the telescope systems were shut down and during the evening when everything was running. Noise estimates were done on both sets of data which confirm that these systems have no net effect on the measure read noise. Night sky data are being used to test the background variance as a function of mean flux. This work is ongoing. Offset pointing: Tests for the accuracy of offsetting were done by a series of observations with increasingly large jitter offsets. The jitter offsets as reported in the headers were compared with the true offsets determined by the stellar positions. The difference between the expected and true offsets showed that the alignment between the detectors and the autoguider CCD was not well determined. We have subsequently been sent raw autoguider CCD images for which we have worked out high quality world coordinate systems. LOWFS world coordinates: Similar work has been done on LOWFS images to determine the world coordinate system for each of these detectors. This work continues. Scattered Light: Observations at varying distances from the full moon have been done to determine whether science observations are likely to be affected badly by scattered light. The analysis of these data continues and we await the arrival of more of the data for this test. Vignetting: A great deal of work has been done to try and work out the vignetting profile of the field of view. Unfortunately it has been impossible to decouple the large scale variation in the detector response from any real vignetting profile in the data. Some of the data for these tests and for further tests which I haven't started yet have just arrived on the new disc so I'll get started on these early next week. MJI said he wanted a series of dark exposures done (5 sec, 10 sec etc) JRL said CRVALs are a serious problem in the recent data and he has spoken to Steven Berad about it. MJI said that the next attempt by ESO to commission VISTA would be early July. A draft report is being prepared for the VISTA project office. JRL will make available all processed data via the webpage in due course. 8. AOB SC talked about Network Area storage options - in particular an Open Solaris system with a Debian front end and a GUI. This involves a licence fee and a maintenance fee with the University possibly paying the latter. Drawbacks - the hardware and software come from different companies. Its probably too expensive. SC has also been looking at UPS facilities for Gaia. Mike reported on recent (and ongoing) problems with the air conditioning in the large computer room. He contacted the workshop who talked to Estate Management who rang the company involved leading to an inordinate delay. Next time we should go straight to the company. A high temperature alarm will be fitted which will be attached to the security alarm. The powers-that-be say we need a fireproof door on the large computer room but we will wait and see on this one.